A set of data contains 57 observations. The lowest value is 41 and the largest is 148. The data are to be organized into a frequency distribution.

a. How many classes would you suggest?



b. What would you suggest as the lower limit of the first class?
1
Expert's answer
2020-02-26T10:58:47-0500

a. If "k" is the number of classes and "n" is the number of observations, then for class intervals, select the smallest "k" such that "2^k>n."

Given that "n=57"


"2^5<57<2^6"

We can take 6 classes.


b. If "U" is the highest observed value and "L" is the lowest observed value, and "w" is the class width (class interval), then we can select "w" such that


"w\geq{U-L \over k}"

Given that "U=148, L=41, k=6." Class limits have the same accuracy as the data values.

Then


"w\geq{148-41 \over 6}""w\geq18"

Use the minimum data entry "L=41" as the lower limit of the first class


"\begin{matrix}
 Class\ limits & Class\ boundaries \\
 41-58 & 40.5-58.5 \\
 59-76 & 58.5-76.5 \\
 77-94 & 76.5-95.5 \\
 95-112 & 94.5-112.5\\
 113-130 & 112.5-130.5 \\
 131-148 & 130.5-148.5
\end{matrix}"